OpenStudy (anonymous):

In how many ways can a committee consisting of 2 faculty members and 5 students be formed if there are 9 faculty members and 14 students eligible to serve on the committee?

OpenStudy (anonymous):

\[{9\choose2}\times{14\choose5}=\frac{9!}{7!2!}\times\frac{14!}{9!5!}\]